OpenStudy (anonymous):

Teesha is in the French club. There are 39 students in the club. The French teacher will pick 5 students at random to guide visiting students from France. What is the probability that Teesha will not be picked as a guide?

OpenStudy (anonymous):

@ganeshie8 I'm having trouble with this one. I'm thinking that the first fraction would be 5/39 because there's 39 total, and five that could possibly be chosen, but I don't know if that's right and I also am not sure what the second fraction is meant to be.

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

So the teacher can select the first student in 38 ways

ganeshie8 (ganeshie8):

that means the probability for not picking Teesha as first student = 38/39, yes ?

OpenStudy (anonymous):

Yes, that was what I thought.

ganeshie8 (ganeshie8):

After that there will be 38 students remaining, the probability for not picking Teesha as second student = 37/38

OpenStudy (anonymous):

THat's exactly what I wrote down on my paper o.o But then I multiplied them and it got a weird answer

ganeshie8 (ganeshie8):

similarly the probality for not picking Teesha as third, fourth or fifth students owuld be : 36/37, 35/36 and 34/35

ganeshie8 (ganeshie8):

multiplying them gives you http://www.wolframalpha.com/input/?i=38%2F39*37%2F38*36%2F37*35%2F36*34%2F35

OpenStudy (anonymous):

Ohhh, so I needed to multiply ALL of them..I got it now.

OpenStudy (anonymous):

Oh, I got that one. Thank you sooo much. This is one I definitely can't get on my own though..

ganeshie8 (ganeshie8):

distribute 1/2

ganeshie8 (ganeshie8):

divide each term by 2

ganeshie8 (ganeshie8):

\[\large \dfrac{1}{2}(-12m+38)\]

ganeshie8 (ganeshie8):

\[\large \dfrac{-12m}{2} + \dfrac{38}{2}\]

OpenStudy (anonymous):

Okay so then I add the fractions?

OpenStudy (anonymous):

oh wait the answer -6m + 19

ganeshie8 (ganeshie8):

Yep!